OpenStudy (anonymous):

Word Problem: alice can row 5 miles down stream in the same time it takes her to row 2 miles upstream. She rows downstream 4 miles/hr faster than she rows upstream. find alicia's rowing rate each way.

OpenStudy (anonymous):

more proof that your math teacher hates you

OpenStudy (anonymous):

think the answer is \[9\tfrac{1}{3}\]

OpenStudy (anonymous):

rate up stream is R - 4 and rate down stream is R + 4. using \[T=\frac{D}{R}\] and the fact that the time going 2 miles downstream is the same as time going 5 up stream gives \[\frac{2}{R+4}=\frac{5}{R-4}\]

OpenStudy (anonymous):

solve this for R via \[2(R-4)=5(R+4)\] \[3R=28\] \[R=\frac{28}{3}\]

OpenStudy (anonymous):

i guess to actually answer you have to add 4 to \[9\tfrac{1}{3}\] so get the rate downstream and subtract 4 from \[9\tfrac{1}{3}\] to get the rate up stream

OpenStudy (anonymous):

i get \[13\tfrac{1}{3} \] going down and \[5\tfrac{1}{3}\] going up

OpenStudy (anonymous):

r*t = d, t = d/r Let r be the rowing rate upstream Solve the following equation for r \[5/(r + 4) = 2/r \]\[\left\{r=\frac{8}{3},4+r=\frac{20}{3},\frac{20}{3}-\frac{8}{3}=4\right\} \]

OpenStudy (anonymous):

oh right. just says 4 miles faster, not current is 4 miles. robtobey is right and i am wrong

OpenStudy (anonymous):

sorry